Question 4 1 pts Suppose a experiment is conducted to study the relationship between the percentage of a certain drug in theQuestion 5 1 pts Using the information in Question 4. Find the least squares estimate of the slope 1 -0.70 O 0.50 O 0.60 0 -0Question 6 1 pts The table below reports the ages (in years) and the number of hours of sleep in one night by seven adults. AQuestion 7 1 pts Using the information in Question 7. The least squares estimate of the slope 8, is about 0 -0.086 0 -0.753 0Question 8 1 pts Using the information in Question 6 and 7. One wishes to test Ho : B1 = 0 vs H, : B170. Which of the followi
Answer #1

Question 4

((N + ΣΧΥ) - ΣΧΣΥ)/((ΝΣ(Χ2) - (ΣΧ)2] (ΝΣ(Y2) – (ΣΥ)2)
r = ( ( 5 * 37 ) - 15 * 10 ) / ( \sqrt{ ( 5 * 55 ) - ( 15 )^{2} )( 5 * 26 ) - ( 10 )^{2} })
r = 0.903

Question 5
ß1 = ( n \Sigma XY - \Sigma X \Sigma Y) / ( n\Sigma X^{2} - (\Sigma X)^{2})
ß1 = ( 5 * 37 - 15 * 10 ) / ( 5 * 55 - ( 15 )^{2})
ß1 = 0.70

Question 6

((N + ΣΧΥ) - ΣΧΣΥ)/((ΝΣ(Χ2) - (ΣΧ)2] (ΝΣ(Y2) – (ΣΥ)2)
r = ( ( 7 * 1916 ) - 337 * 44 ) / ( \sqrt{ ( 7 * 18563 ) - ( 337 )^{2} )( 7 * 296 ) - ( 44 )^{2} })
r = - 0.95

Question 7

ß1 = ( n \Sigma XY - \Sigma X \Sigma Y) / ( n\Sigma X^{2} - (\Sigma X)^{2})
ß1 = ( 7 * 1916 - 337 * 44 ) / ( 7 * 18563 - ( 337 )^{2})
ß1 = - 0.086

Question 8